OpenStudy (anonymous):

Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8

jimthompson5910 (jim_thompson5910):

If f(x) = 20, then f(x) = 2(x + 4) 20 = 2(x + 4) So you need to solve 20 = 2(x + 4) to find your answer

OpenStudy (anonymous):

what is the answer

jimthompson5910 (jim_thompson5910):

Do you know how to solve the last equation?

OpenStudy (anonymous):

Jim Thompson is right. 20 = 2(x + 4) 20 = 2x + 8 12 = 2x 6 = x The answer is 6. :)

OpenStudy (anonymous):

thankyou !! <3

OpenStudy (anonymous):

You are welcome. c: